# Planet formation regulated by galactic-scale interstellar turbulence

### v1

@@ 0.10 Disc properties (step 6):
To evolve the disc mass, in our fiducial model we fix the initial $M_{\mathrm{disc}}(t=0)=0$. This is equivalent to the assumption that the initial disc that forms due to core collapse is low mass, or rapidly depleted. This runs counter to the conventional assumption that the protoplanetary disc represents an evolved state of the primordial protostellar disc, which formed during the stellar core collapse. We make this decision as a ‘clean’ test of our model; we wish to explore the properties of discs arising from BHL accretion alone. Such a model requires no additional free parameters. We emphasise that no specific observational constraints contradict a low initial disc mass. Any such evidence would need to connect the primordial circumstellar material, which forms over $\sim 10^{2}$ years, to the protoplanetary disc, which in the context of this work forms over $\sim 10^{5}$ years and evolves over $\sim 10^{6}$ years. Although (massive) discs around Class 0 stars have been inferred,[27, 91] interpreting these observations in terms of the exact age of the (proto-)star (down to $\sim 10^{4}$ years), the nature of the disc and the time-scale over which any such disc should survive is not simple. Since few, if any, unambiguous observations exist constraining the properties of primordial discs that are categorically around stars of age $\ll 10^{5}$ years, there is no reason to assume that such a primordial disc should be both long-lived and comparable in mass to the subsequently accreted one. Indeed, it would be a remarkable coincidence if BHL accretion and protostar formation contribute an equal fraction of mass to the disc population. Nonetheless, we also explore the impact of this assumption on our results below.
@@
For $\tau_{\mathrm{acc}}=M_{\mathrm{disc}}/\dot{M}_{\mathrm{acc}}$ we assume a log-uniform distribution ±$\pm 1$ dex around $\log\tau_{\mathrm{acc}}/\mathrm{1\,Myr}=-0.5$, based on visual inspection of the observed distribution.[44] Empirically, the time-scale $\tau_{\mathrm{acc}}$ is not strongly correlated with stellar mass, but may be a complex function of the disc properties or stellar age. For example, visually interpreting Figure 12 of Almendros-Abad et al. (2024),[44] the median accretion time-scale appears to be $\tau_{\mathrm{acc,1/2}}\sim 0.1$ Myr for the young region $\rho$ Ophiuchus, while $\tau_{\mathrm{acc,1/2}}\sim 1$ Myr for the older Lupus region. Indeed, we expect rapid BHL accretion to instigate turbulence in the disc, temporarily decreasing $\tau_{\mathrm{acc}}$. However, empirically it is not clear the degree to which any variation of $\tau_{\mathrm{acc}}$ is a survival bias, discs with shorter $\tau_{\mathrm{acc}}$ having been depleted (we will show a similar evolution in our model which has fixed $\tau_{\mathrm{acc}}$ for each star below). While exploring time- or accretion-dependent $\tau_{\mathrm{acc}}$ is an interesting avenue for future exploration, we initially maintain our simplified model, subsequently exploring the potential role of the ISM in driving a time-dependent stellar accretion rate.
@@
0.11 Observational comparison (step 7):
We wish to compare our model to observed discs with comparable stellar ages. The ages of star forming regions can both be observationally uncertain[92] and should exhibit physical scatter given the finite duration of the star formation process. However, nominally ages for the youngest populations of stars in the $\rho$ Ophiuchus complex are $\sim 0.1-1$ Myr,[93] in Taurus are $\sim 1{-}3$ Myr,[94] in Lupus are $\lesssim 3$ Myr,[95] in Chameleon I have been estimated at $\sim 1{-}2$ Myr[96] and $\sim 3{-}6$ Myr,[97] and Upper Sco are $\sim 5-10$ Myr.[98, 99] We compare these regions to our model at ages, $0.5$, $1$, $2$, $3$ and $5$ Myr respectively; these choices should be considered illustrative, not precise. For total disc masses we take $M_{\mathrm{disc}}=10^{2}M_{\mathrm{dust}}$, where $M_{\mathrm{dust}}$ is as quoted by Manara et al. (2023),[42] inferred from the mm continuum flux with standard assumptions for opacity and temperature. In order to draw comparisons with each, we perform a Gaussian kernel density estimate with Scott’s rule for bandwidth selection.[100] Where we compare to data directly, we weight the model data points to yield the same distribution in the stellar mass axis as for the observational dataset. In practice, this means dividing the stellar mass distribution into bins, and defining the weight of model data points in each bin $i$ to be $N_{\mathrm{obs},i}/N_{\mathrm{mod},i}$, where $N_{\mathrm{obs},i}$ and $N_{\mathrm{model},i}$ are the number of observed and model points in bin $i$ respectively. The results of these comparisons are shown in Figure 2 and discussed in the main article, however we make some additional comments on the disc properties here as follows.
0.12 Disc masses:
@@
We show the time evolution of the disc mass distribution in our model more quantitatively in Extended Data Figure 3, where we plot the cumulative distribution function for disc masses around stars $0.5\,M_{\odot}<m_{*}<1\,M_{\odot}$. We also show observed distributions for Taurus and Upper Sco, leaving out the other regions for clarity. We can see that at $\sim 1$ Myr, the disc mass distribution in our model is very similar to the distribution in Taurus. This decreases after a few Myr until reaching a similar distribution to that seen in Upper Sco between $3$ and $9$ Myr. Within uncertainties, and given the caveat of a range of stellar ages in each region, our model appears fully consistent with the observed disc mass distributions. More speculatively, we note that the disc takes $\sim 0.3$ Myr to reach the maximum mass distribution. This is interesting in the context of $\rho$ Ophiuchus, which have discs that are more compact[45, 46] and lower mass[101, 47] than $\sim 1{-}2$ Myr old regions. A similar result has been found in Corona Australis.[102] Our models suggest that this is plausibly the result of ongoing mass assembly proceeding from outside-in, via BHL accretion.
0.13 Stellar accretion rates:
@@
The accretion rates from our model are entirely determined by the disc mass distribution, since we have imposed $\dot{M}_{\mathrm{acc}}\propto M_{\mathrm{disc}}/\tau_{\mathrm{acc}}$. They are therefore not an independent test of our model, but a validation or repudiation of this assumption. Given that we have anchored $\tau_{\mathrm{acc}}$ to observations, and we have found observational agreement with the disc mass distribution, as expected we also find for the majority of regions we have excellent agreement for accretion rates (middle row, Figure 2). We note minor tension particularly in the case of Chameleon I, for which some accretion rates are somewhat larger than predicted by our model. This is plausibly the result of a younger population than the comparison age we have adopted. For example, Galli et al. (2021)[96] argue that the older esimate obtained by Luhman et al. (2007)[97] may be explained by systematic distance underestimate (then Chameleon I may be $\sim 1{-2}$ Myr old). Overall, we find excellent agreement between stellar accretion rates in our model and the observed samples. Since our model does not contain any stellar mass dependent disc physics (except the BHL accretion rate), this makes a strong argument in favour of BHL accretion as the origin of disc assembly.
@@
We may also ask whether the data supports our assumption of a fixed $\tau_{\mathrm{acc}}$ distribution with stellar mass and time. It is possible, even probable, that this accretion time-scale does vary over time, particularly if internal disc turbulence is driven to some degree by the material inherited from the ISM (which is plausible, see Figure 4). As previously discussed, empirically there is some variation in $\tau_{\mathrm{acc}}$ between regions.[44] In this case, the true distribution of $\tau_{\mathrm{acc}}$ may be a function of the state of the ISM in each specific region. However, even for fixed $\tau_{\mathrm{acc}}$ for individual stars, the fact that star-disc systems with shorter $\tau_{\mathrm{acc}}$ lose their discs earlier mean that they drop out of the sample of surviving discs. For our model, this is shown in Extended Data Figure 4, where we find that once discs are $\sim 3$ Myr old, the median $\tau_{\mathrm{acc}}\sim 1$ Myr, a factor three larger than for the initial disc population. This moderate increase is similar to the median for the observed distribution for discs of this age.[44] We conclude that, while $\tau_{\mathrm{acc}}$ may vary over time for individual stars, such temporal variation is not required to produce the observed disc population.
0.14 Outer radii:
@@
We can also consider the distribution of gaseous disc radii $R_{\mathrm{disc}}$ we infer from our model. However, the appropriate definition of $R_{\mathrm{disc}}$ is less clear than for disc masses or stellar accretion rates in terms of the correspondence between the model and observational data. The observed outer radius in dust is dependent on variations in the dust opacity [103] and in gas molecular tracers, such as CO, outer radii are dependent on thermochemistry.[104] Both are dependent on the definition of ‘outer radius’, often defined by the radius that encloses some fraction of the total flux. In our simple models, we do not explicitly define an outer radius. A sensible estimate for $R_{\mathrm{disc}}$ is the accretion radius $R_{\mathrm{acc}}$ at the most recent time for which the disc is accumulating mass (i.e. $\dot{M}_{\mathrm{BHL}}>\dot{M}_{\mathrm{acc}}$), and this is the definition we adopt in this work and in constructing Figure 2. In the absence of other disc evolution processes, such as viscous spreading,[89] we actually expect $R_{\mathrm{acc}}>R_{\mathrm{disc}}$, since the outer radius depends on the fraction of residual angular momentum from the accreted ISM. With these caveats, we proceed with the assumption $R_{\mathrm{acc}}\approx R_{\mathrm{disc}}$ in our comparison with observations in the main article.
0.15 Disc lifetimes:
@@
For a fair comparison of disc property distributions, we must also confirm that the total fraction of discs included in our sample is similar in our model to those in star forming regions of a given age. We therefore show the disc fraction as a function of time in Extended Data Figure 2, by the definition $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$. We confirm that our model yields reasonable disc fractions over the range of times we consider in our disc property comparisons. We find a very similar disc fraction evolution to the exponential decay relationships suggested in the literature, typically with dispersal time-scale $\tau_{\mathrm{disc}}\sim 3{-}5$ Myr. [1, 105]
@@
Interestingly, we find that the initial disc fraction is not $100$ percent in our model, but is maximal at $\sim 85$ percent. This appears to be consistent with observations which indicate that, even for very young disc populations, the disc fraction never exceeds $\sim 80$ percent.[106] However, when accounting for close binaries, this fraction could be closer to $100$ percent for single stars.[107] The model prediction of a somewhat suppressed initial disc fraction should therefore not be interpreted as a unique signature of BHL accretion, and may also depend on the presence or absence of a primordial disc. We also find a somewhat flatter disc fraction as a function of time than a simple exponential decay with linear time exponent. Some authors to suggest such a flattened functional form for the disc fraction,[78] supported by a non-negligible disc fractions in clusters that are tens of Myr old.[108] Such surviving discs can be easily understood in the context of the stochastic accretion process that defines our model.
@@
We caveat this discussion in that we have not included any end-stage disc dispersal processes; in particular the photoevaporative wind driven by the central star.[109, 110, 111, 112, 113] We would expect that, at least for very low mass discs, mass-loss may be dominated by such a wind. Our assumption is effectively that mass-loss in the wind is inefficient; in this case we can assume that the lifetime of an observable disc is essentially set by the accretion process in isolation. Once a disc is ‘dispersed’ in our model, if at some later stage the disc replenishment occurs more rapidly than the true photoevaporation rate, then our model remains valid. Nonetheless, a future investigation of how photoevaporation may interact with BHL accretion for disc populations is warranted in future work.

### v2

@@ 3.1 Simple disc model
To evolve the disc mass, in our fiducial model we fix the initial $M_{\mathrm{disc}}(t=0)=0$. This is equivalent to the assumption that the initial disc that forms due to core collapse is low mass, or rapidly depleted. This runs counter to the conventional assumption that the protoplanetary disc represents an evolved state of the primordial protostellar disc, which formed during the stellar core collapse. We make this decision as a ‘clean’ test of our model; we wish to explore the properties of discs arising from BHL accretion alone. Such a model requires no additional free parameters. We emphasise that no specific observational constraints contradict a low initial disc mass. Any such evidence would need to connect the primordial circumstellar material, which forms over $\sim 10^{2}$ years, to the protoplanetary disc, which in the context of this work forms over $\sim 10^{5}$ years and evolves over $\sim 10^{6}$ years. Although (massive) discs around Class 0 stars have been inferred (Maury et al., 2019; Tobin et al., 2020) interpreting these observations in terms of the exact age of the (proto-)star (down to $\sim 10^{4}$ years), the nature of the disc and the time-scale over which any such disc should survive is not simple. Since few, if any, unambiguous observations exist constraining the properties of primordial discs that are categorically around stars of age $\ll 10^{5}$ years, there is no reason to assume that such a primordial disc should be both long-lived and comparable in mass to the subsequently accreted one. Indeed, it would be a remarkable coincidence if BHL accretion and protostar formation contribute an equal fraction of mass to the disc population. Nonetheless, we also explore the impact of this assumption on our results below.
@@
For $\tau_{\mathrm{acc}}=M_{\mathrm{disc}}/\dot{M}_{\mathrm{acc}}$ we assume a log-uniform distribution ±$\pm 1$ dex around $\log\tau_{\mathrm{acc}}/\mathrm{1\,Myr}=-0.5$, based on visual inspection of the observed distribution (Almendros-Abad et al., 2024). Empirically, the time-scale $\tau_{\mathrm{acc}}$ is not strongly correlated with stellar mass, but may be a complex function of the disc properties or stellar age. For example, visually interpreting Figure 12 of Almendros-Abad et al. (2024), the median accretion time-scale appears to be $\tau_{\mathrm{acc,1/2}}\sim 0.1$ Myr for the young region $\rho$ Ophiuchus, while $\tau_{\mathrm{acc,1/2}}\sim 1$ Myr for the older Lupus region. Indeed, we expect rapid BHL accretion to instigate turbulence in the disc, temporarily decreasing $\tau_{\mathrm{acc}}$. However, empirically it is not clear the degree to which any variation of $\tau_{\mathrm{acc}}$ is a survival bias, discs with shorter $\tau_{\mathrm{acc}}$ having been depleted (we will show a similar evolution in our model which has fixed $\tau_{\mathrm{acc}}$ for each star below). While exploring time- or accretion-dependent $\tau_{\mathrm{acc}}$ is an interesting avenue for future exploration, we initially maintain our simplified model, subsequently exploring the potential role of the ISM in driving a time-dependent stellar accretion rate.
@@
3.2 Observational comparison
We wish to compare our model to observed discs with comparable stellar ages. The ages of star forming regions can both be observationally uncertain (e.g. Soderblom et al., 2014) and should exhibit physical scatter given the finite duration of the star formation process. However, nominally ages for the youngest populations of stars in the $\rho$ Ophiuchus complex are $\sim 0.1-1$ Myr (Luhman & Rieke, 1999) in Taurus are $\sim 1{-}3$ Myr (Krolikowski et al., 2021), in Lupus are $\lesssim 3$ Myr (Galli et al., 2020) in Chameleon I have been estimated at $\sim 1{-}2$ Myr (Galli et al., 2021b) and $\sim 3{-}6$ Myr (Luhman, 2007), and Upper Sco are $\sim 5-10$ Myr (Pecaut & Mamajek, 2016; Luhman & Esplin, 2020). We compare these regions to our model at ages, $0.5$, $1$, $2$, $3$ and $5$ Myr respectively; these choices should be considered illustrative, not precise. For total disc masses we take $M_{\mathrm{disc}}=10^{2}M_{\mathrm{dust}}$, where $M_{\mathrm{dust}}$ is as quoted by Manara et al. (2023), inferred from the mm continuum flux with standard assumptions for opacity and temperature. In order to draw comparisons with each, we perform a Gaussian kernel density estimate with Scott’s rule for bandwidth selection (Scott, 1992). Where we compare to data directly, we weight the model data points to yield the same distribution in the stellar mass axis as for the observational dataset. In practice, this means dividing the stellar mass distribution into bins, and defining the weight of model data points in each bin $i$ to be $N_{\mathrm{obs},i}/N_{\mathrm{mod},i}$, where $N_{\mathrm{obs},i}$ and $N_{\mathrm{model},i}$ are the number of observed and model points in bin $i$ respectively. The results of these comparisons are shown in Figure 3, in which we find remarkable agreement with the observed disc mass, stellar accretion rate, and outer disc radius distributions. We explore these findings in greater detail as follows.
Figure 3: Linear kernel density estimates (KDEs) for the disc mass (top), stellar accretion rate (middle) and accretion radii (bottom) versus stellar mass in our model compared to observations (red points, red dashed lines) in local star forming regions. We show results at times $0.5$, $1$, $2$ and $3$ and $5$ Myr. The red points represent the distribution of 100×100$\times$100 × dust disc masses (top), stellar accretion rate (middle) in $\rho$ Ophiuchus, Taurus, Lupus, Chameleon I and Upper Scorpius (left to right – Manara et al., 2023). When constructing the KDEs, model points are weighted to yield the same distribution along the stellar mass axis as in the observational samples, except in the bottom row. For comparison with the accretion radii, we show the empirically inferred relationship between CO outer disc radii as a red dashed line in the bottom panels (Andrews, 2020).
3.2.1 Disc masses
@@
We show the time evolution of the disc mass distribution in our model more quantitatively in Figure 4, where we plot the cumulative distribution function for disc masses around stars $0.5\,M_{\odot}<m_{*}<1\,M_{\odot}$. We also show observed distributions for Taurus and Upper Sco, leaving out the other regions for clarity. We can see that at $\sim 1$ Myr, the disc mass distribution in our model is very similar to the distribution in Taurus. This decreases after a few Myr until reaching a similar distribution to that seen in Upper Sco between $3$ and $9$ Myr. Within uncertainties, and given the caveat of a range of stellar ages in each region, our model appears fully consistent with the observed disc mass distributions. More speculatively, we note that the disc takes $\sim 0.3$ Myr to reach the maximum mass distribution. This is interesting in the context of $\rho$ Ophiuchus, which have discs that are more compact (Testi et al., 2016; Cieza et al., 2019) and lower mass (Williams et al., 2019; Testi et al., 2022) than $\sim 1{-}2$ Myr old regions. A similar result has been found in Corona Australis (Cazzoletti et al., 2019). Our models suggest that this is plausibly the result of ongoing mass assembly proceeding from outside-in, via BHL accretion.
Figure 4: Cumulative distribution of disc masses at different times in our model (solid lines) compared to the observed distribution (dashed lines) for Taurus (blue) and Upper Sco (orange). We include only two observed disc populations for clarify. Shaded region represents the statistical uncertainty interval for observational datasets.
3.2.2 Stellar accretion rates
@@
The accretion rates from our model are entirely determined by the disc mass distribution, since we have imposed $\dot{M}_{\mathrm{acc}}\propto M_{\mathrm{disc}}/\tau_{\mathrm{acc}}$. They are therefore not an independent test of our model, but a validation or repudiation of this assumption. Given that we have anchored $\tau_{\mathrm{acc}}$ to observations, and we have found observational agreement with the disc mass distribution, as expected we also find for the majority of regions we have excellent agreement for accretion rates (middle row, Figure 3). We note minor tension particularly in the case of Chameleon I, for which some accretion rates are somewhat larger than predicted by our model. This is plausibly the result of a younger population than the comparison age we have adopted. For example, Galli et al. (2021b) argue that the older esimate obtained by Luhman (2007) may be explained by systematic distance underestimate (then Chameleon I may be $\sim 1{-2}$ Myr old). Overall, we find excellent agreement between stellar accretion rates in our model and the observed samples. Since our model does not contain any stellar mass dependent disc physics (except the BHL accretion rate), this makes a strong argument in favour of BHL accretion as the origin of disc assembly.
Figure 5: Cumulative distribution of the accretion time-scales $\tau_{\mathrm{acc}}$ at different times for surviving discs in our model. This evolution is purely due to discs with short accretion time-scales dropping out of the sample.
@@
We may also ask whether the data supports our assumption of a fixed $\tau_{\mathrm{acc}}$ distribution with stellar mass and time. It is possible, even probable, that this accretion time-scale does vary over time, particularly if internal disc turbulence is driven to some degree by the material inherited from the ISM (which is plausible, see Figure 11). As previously discussed, empirically there is some variation in $\tau_{\mathrm{acc}}$ between regions (Almendros-Abad et al., 2024). In this case, the true distribution of $\tau_{\mathrm{acc}}$ may be a function of the state of the ISM in each specific region. However, even for fixed $\tau_{\mathrm{acc}}$ for individual stars, the fact that star-disc systems with shorter $\tau_{\mathrm{acc}}$ lose their discs earlier mean that they drop out of the sample of surviving discs. For our model, this is shown in Figure 5, where we find that once discs are $\sim 3$ Myr old, the median $\tau_{\mathrm{acc}}\sim 1$ Myr, a factor three larger than for the initial disc population. This moderate increase is similar to the median for the observed distribution for discs of this age (Almendros-Abad et al., 2024). We conclude that, while $\tau_{\mathrm{acc}}$ may vary over time for individual stars, such temporal variation is not required to produce the observed disc population.
3.2.3 Outer disc radii
@@
We can also consider the distribution of gaseous disc radii $R_{\mathrm{disc}}$ we infer from our model. However, the appropriate definition of $R_{\mathrm{disc}}$ is less clear than for disc masses or stellar accretion rates in terms of the correspondence between the model and observational data. The observed outer radius in dust is dependent on variations in the dust opacity (Rosotti et al., 2019) and in gas molecular tracers, such as CO, outer radii are dependent on thermochemistry (Trapman et al., 2023). Both are dependent on the definition of ‘outer radius’, often defined by the radius that encloses some fraction of the total flux. In our simple models, we do not explicitly define an outer radius. A sensible estimate for $R_{\mathrm{disc}}$ is the accretion radius $R_{\mathrm{acc}}$ at the most recent time for which the disc is accumulating mass (i.e. $\dot{M}_{\mathrm{BHL}}>\dot{M}_{\mathrm{acc}}$), and this is the definition we adopt in this work and in constructing Figure 3. In the absence of other disc evolution processes, such as viscous spreading (Lynden-Bell & Pringle, 1974), we actually expect $R_{\mathrm{acc}}>R_{\mathrm{disc}}$, since the outer radius depends on the fraction of residual angular momentum from the accreted ISM. With these caveats, we proceed with the assumption $R_{\mathrm{acc}}\approx R_{\mathrm{disc}}$ in our comparison with observations in the main article.
3.2.4 Disc lifetimes
@@
For a fair comparison of disc property distributions, we must also confirm that the total fraction of discs included in our sample is similar in our model to those in star forming regions of a given age. We therefore show the disc fraction as a function of time in Figure 6, by the definition $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$. We confirm that our model yields reasonable disc fractions over the range of times we consider in our disc property comparisons. We find a very similar disc fraction evolution to the exponential decay relationships suggested in the literature, typically with dispersal time-scale $\tau_{\mathrm{disc}}\sim 3{-}5$ Myr (Haisch et al., 2001; Ribas et al., 2015).
Figure 6: Fraction of surviving discs in our model (black line), defined to be discs which have a mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ for stars with masses in the range $0.1{-}3\,M_{\odot}$. We show the evolution of disc fraction given by $\exp\left(-t/\tau_{\mathrm{disc}}\right)$ for $\tau_{\mathrm{disc}}=3$ Myr (Haisch et al., 2001) and $5$ Myr (Ribas et al., 2015) as blue and red lines respectively.
@@
Interestingly, we find that the initial disc fraction is not $100$ percent in our model, but is maximal at $\sim 90$ percent. Observations indicate that, even for very young disc populations, the disc fraction never exceeds $\sim 80$ percent (Michel et al., 2021). However, when accounting for close binaries, this fraction could be closer to $100$ percent for single stars (Kraus et al., 2012). The model prediction of a somewhat suppressed initial disc fraction should therefore not be interpreted as a unique signature of BHL accretion, and may also depend on the presence or absence of a primordial disc. We also find a somewhat flatter disc fraction as a function of time than a simple exponential decay with linear time exponent. Some authors to suggest such a flattened functional form for the disc fraction (Pfalzner et al., 2022), supported by a non-negligible disc fractions in clusters that are tens of Myr old (Galli et al., 2021a). Such surviving discs can be easily understood in the context of the stochastic accretion process that defines our model.
@@
We caveat this discussion in that we have not included any end-stage disc dispersal processes; in particular the photoevaporative wind driven by the central star (Clarke et al., 2001; Alexander et al., 2006; Gorti & Hollenbach, 2009; Owen et al., 2010; Picogna et al., 2019). We would expect that, at least for very low mass discs, mass-loss may be dominated by such a wind. Our assumption is effectively that mass-loss in the wind is inefficient; in this case we can assume that the lifetime of an observable disc is essentially set by the accretion process in isolation. Once a disc is ‘dispersed’ in our model, if at some later stage the disc replenishment occurs more rapidly than the true photoevaporation rate, then our model remains valid. Nonetheless, a future investigation of how photoevaporation may interact with BHL accretion for disc populations is warranted in future work.